Astronomy Report – February 06th

Astronomy Report

Name: Ghanim Alotaibi

Crew: 221
Date: 06 Feb 2020

A mistake was found in the coordinate submitted for the patterns of the discovery project. The discovery patterns observations submitted yesterday was therefore canceled from Skynet. New patterns will be identified for tomorrow’s night and will be submitted to Skynet.

A new variable star target was identified. The astronomy proposal was updated and sent to observatory director.


Robotic Telescope Requested MDRS-14

Objects to be Imaged this Evening:

  1. V CAS Variable
  2. 2MASS J06593158-0405277(new target submitted today)
  3. Orion Nebula

Images submitted with this report: Non

Problems Encountered: Non



Operations Report – February 06th

Crew 221 Operations Report 06Feb2022

SOL: 4

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 8pm)

Hours run: 13:00

From what time last night: 7:00pm

To what time this morning: 8:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 76%

Diesel Reading – 98%

Station Propane Reading – 61%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 355 gals

Static to Loft Pump used – yes

Water in GreenHab – 243 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 35%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 135

Beginning charge: 100%

Ending charge: 25%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: not used

Summary of Hab operations: nominal.

Summary of GreenHab operations: black corn experiment started, 4 trays were filled 3 with mars soil and one with regular soil for control, beans were plantes and watered with 2L. operational, watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators runing, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly is ongoing adn expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: no issues

Questions, concerns and requests to Mission Support:

GreenHab Report – February 06th

Crew 221 GreenHab Report 06FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average temperatures: 18.33°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 9.2 gallons

Daily water usage for research and/or other purposes: 0.5 gallons

Water in Blue Tank: 243 gallons

Time(s) of watering for crops: 1

Changes to crops: Two of the four succulent plants seemed to have recovered a bit, yay! The leaves are more plump. Waiting on the other two. I repotted one of the sadder looking ones, as the soil was completely dry and barren and the roots didn’t seem healthy. Hanging lettuce pot on the bottom was fully soaked in water. As the pot has no drainage, will leave lettuce unwatered for the next day in the hopes that the soil dries out a bit. A few tomato plants are wilting. Rotated the plants on the bed. Ginger has almost sprouted, and the radish sprinklers, bell peppers and roma tomato have all sprouted.

Narrative: The plant beds were very moist this morning so I skipped the morning watering. Crew Engineer and I set up a new experiment in the GreenHab. We are going to attempt to grow Blue Corn from Martian soil! We have set up four trays on one of the benches, one as the control and three containing different types of Martian soil, mixed with some fertilizer. We’ve planted several rows of blue corn and will be watering it and monitoring the soil conditions each day.

Harvest: NA

Support/supplies needed: NA

Other Notes: Have sprayed the herb garden and the carrots and icy radishes with some pesticide to get rid of the small bugs. Will not harvest for the next few days. Also, slightly concerned about the wilty tomato plants. I have no experience with tomatos. Would this be a symptom of too much/little water? Or too much/little sunlight?

Sol Summary – February 06th


Summary Title: It’s all about the Science

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol Summary: Crew 221 had great Mexican tacos and prepared a new batch of kombucha that they inherited from crew 220. It will be ready by the 11thof February! No big evening activities were planned as the crew needed some good night sleep to refill their batteries. The discussions between the crew members, however, become more and more interesting each day.

This morning, the commander woke up the team with “starships” by Nicki Minaj to get the team up and running and also prepared pancakes with fresh banana for the crew.

After a short team meeting to plan the day, the crew teamed up in pairs. Crew Engineer L. Monge and GreenHab Officer J.Todd focused on ISRU plant experiments in the GreenHab to plant black corn in different soils taken during yesterday’s EVA from different places. Crew Journalist H. Kalucha and Crew Astronomer G. Alotaibi started to build the crew robot in the RAM that shall be tested once ready during EVAs, even though several crew members dropped by to help as everyone was excited to build it. Crew Commander M.Grulich and HSO R. Alshammari were preparing the soil samples from the last EVA and inspecting the petri dishes from yesterday. They were positively surprised to find the first signs of bacteria growth from the sample of the EVA space suit before it was cleaned. The air sample and soil sample from yesterday did not show any sign of bacteria growth yet.

In the evening, the crew will have Costa-Rican empanadas and another meditation session shortly before the Australian cultural night, where J. Todd will try to sell the crew on vegemite.

EVA: None

Reports to be filed:

Sol Summary

Journalist Report

Photo Report

EVA Request

Operations Report

GreenHab Report

Astronomy Report

Science Report

Support Requested: None

Journalist Report – February 06th

Crew 221 Journalist Report 06Feb2020

Author: Hemani Kalucha, Crew Journalist

Sol 4

Title: A day of science!

This morning, we were awoken by our wonderful commander making pancakes and serving the last of the fresh fruit – bananas. Energized by this meal, the crew divided into teams of three. Luis and Jess planted the black corn seeds in the greenhab using martian soil, collected from the EVA done yesterday. Maria and Rawan spent the whole day growing bacteria in the samples also collected from the EVA done yesterday. One of the samples saw signs of bacteria, which caused lots of excitement and a feeling of success amongst the crew. Hemani and Ghanim started to build the robot. After a fun team lunch of good conversation and two-minute Ramen, the teams switched around, and the all girl engineering team (Hemani, Maria, Jess) put together the robotic arm! It started to get cold in the RAM, so the Crew Engineer didn’t have time to get his moisture sensor fully working, but all in all, much progress was made on all fronts. The mini missions made the hours go by a little faster, since the lack of internet and space, usually makes the day feel long and slow. Tonight will be another fun filled evening with sports, meditation, and another crewmember’s cultural exploration.

Science Report – February 06th

Science Report

Crew: 221

Date: 06-FEB-20

Crew Science Officer: Hemani Kalucha

1. Microbiology Research Today, we experimented with the microbiology samples collected from yesterday’s EVA. Each sample was weighed and one gram of each sample was taken for futher preparation. The 1 gram was put into 100ml of distilled water and then diluted further into 9ml of distilled water with 1 ml of of the previous solution (10^-3 concentration of potential bacteria). A small amount of the final solution was then spread onto the petri dishes to be incubated at 25degC. The samples from yesterday were inspected and the first signs of bacteria growth was identified on the sample of the space suit before we cleaned the helmet. This sample will be further analysed in the coming days. A first attempt to prepare our own agar was done but resulted in a spill which was cleaned and a new attempt will be done tomorrow.

2. ISRU project Today we set up four trays – one control fertilized soil with perlite and three different types of mars regolith, collected on EVA 4. Within each sample, half was fertilized and half was not. Then, several rows of black corn were inserted and kept in direct sunlight in the greenhab. Each tray was watered with 0.5 litres. The soil health and crop health will be monitored over the next week.

Copyright © The Mars Society. All rights reserved. | Main Site